OpenStudy (anonymous):

What is 3 root of − 320 end root? A. − 5 3 root of 4 B. 4 3 root of − 4 end root C. − 4 3 root of 5 D. 5 3 root of 6

OpenStudy (michele_laino):

here, we have: \[\Large \sqrt[3]{{ - 320}} = \sqrt[3]{{ - {4^3} \times 5}}\]

OpenStudy (anonymous):

i-think-its-c

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay